OpenStudy (anonymous):

Lim as x approaches sin(2x) / 4x?

OpenStudy (anonymous):

pls help @lgbassalote @kevinkeegan

OpenStudy (mathmagician):

if x approaches to 0, you know that sin(x)/x=1 sosin(2x)/4x=sin(2x)/(2*2x))=1/2